# Imperfect Competition in Markets for Short-Circuit Current Services

### v2

@@ 3.2 Primal-Dual Single-Level Optimization Model
It is noteworthy that problem (7) is still nonlinear, as it still contains quadratic terms $\lambda^{\textrm{E}}_{t}P_{\hat{g},t}$, $\lambda_{b,t}^{\textrm{SCC}}\textrm{k}_{b\hat{g}}u_{\hat{g},t}$ in (3a), $\textrm{O}_{\hat{g},b,t}^{\textrm{SCC}}\beta_{\hat{g},b,t}^{\textrm{SCC}}u_{\hat{g},t}$, $\textrm{O}_{\hat{g}}^{\textrm{m}}\beta_{\hat{g},t}^{\textrm{m}}P_{\hat{g},t}$ in (4a). McCormick envelopes are employed to perform the linearization with proper upper bounds on relevant variables, as the setting in [25]. The envelopes for those second-order terms are shown as follows.
@@
In summary, to solve the nonlinear bilevel model with non-convexity, binary variables are relaxed to be continuous to derive the dual of primal LL problem. Then they are tightened to achieve realistic UC solutions, although this leads to a DG between the primal and dual LL problems. The penalty function method is used to minimize the DG, giving a trade-off between the solution accuracy and values of objectives. Finally, nonlinear terms are restated to form the final single-level, mixed-integer linear program.
4 Case Studies
4.1 Test System Setting
Figure 1: Modified IEEE 30-bus system.
Table 1: Operating Parameters of Synchronous Generators
Bus 2 3 4 5 27 30
$\textrm{O}_{g}^{\textrm{nl}}$ (€/h) 1,743 1,501 1,376 1,093 990 857
$\textrm{O}_{g_{1}}^{\textrm{m}}$ (€/MWh) 6.20 7.10 10.47 12.28 13.53 15.36
$\textrm{O}_{g_{2}}^{\textrm{m}}$ (€/MWh) 7.07 8.72 11.49 12.84 14.60 15.02
$\textrm{K}_{g}^{\textrm{st}}$ (€/h) 20,000 12,500 9,250 7,200 5,500 3,100
$\textrm{K}_{g}^{\textrm{sh}}$ (€/h) 5,000 2,850 1,850 1,440 1,200 1,000
$\textrm{P}_{g}^{\textrm{min}}$ (MW) 658 576 302 133 130 58
$\textrm{P}_{g}^{\textrm{max}}$ (MW) 1,317 1,152 756 667 650 576
$\textrm{P}_{g}^{\textrm{rd}}$ (MW/h) 263 230 302 267 390 346
$\textrm{P}_{g}^{\textrm{ru}}$ (MW/h) 263 230 302 267 390 346
$u_{g,0}$ 1 1 1 1 1 0
$P_{g,0}$ (MW) 1,054 922 605 534 520 0
Case studies on a modified IEEE 30-bus system are carried out in this work. The diagram of the test system is shown in Fig. 1. The IBR, i.e., wind turbines are added to buses {1, 23, 26}, with energy bids {2.46, 3.19, 2.86} €/MWh, respectively. In order to model a modular SCC contribution from SGs, we consider that two units are located in each bus containing an SG, i.e., buses {2, 3, 4, 5, 27, 30}. This way, diverse decisions can be made in the market. For example, different UC status (i.e., both online, both offline, just one of them online) can determine various SCC contributions and bids. The SGs in bus 2 are separately named as $g_{1}$-bb2 and $g_{2}$-bb2, with 2gg-bb2 standing for both of them. Other thermal units follow the same nomenclature. The SGs’ parameters are given in Table 1.
The network impedances and reactances of SGs refer to [28] and [29], respectively. Load demand $\textrm{P}^{\textrm{D}}_{t}\in[\textrm{5109},\textrm{7689}]$ MW. Total installed wind generation capacity is 750 MW. Base power is $S_{B}=\textrm{100}~\textrm{MVA}$. $\textrm{I}_{c}$, the short-circuit injection from IBR, is 1.00 p.u. Nominal voltage is set as 0.95 p.u. for conservativeness [8]. Maximum multipliers for bidding decisions $\overline{\upbeta}^{\textrm{m}}_{\hat{g}}$ and $\overline{\upbeta}^{\textrm{SCC}}_{\hat{g},b}$ are both set as 2. All code and data are available at the public repository [26]. The simulations in this work were processed by Julia-JuMP and Gurobi in version 12.0.0.
4.2 Validation of Short-Circuit Current Constraints
The approximate SCC constraints (2a) are linearized to be (4c) for the primal-dual formulation. Two evaluations are employed to validate those SCC constraints in the first place.
Evaluation 1: Training errors of approximation. Two types of errors are analyzed to verify the approximation [7]: Type-I: $I_{b_{\textrm{L}}}\geq\textrm{I}_{b_{\textrm{lim}}}\cap I_{{b_{\textrm{SC}}}}<\textrm{I}_{b_{\textrm{lim}}}$ means the approximate SCC is looser than the actual requirement, a potentially dangerous situation; Type-II: $I_{b_{\textrm{L}}}<\textrm{I}_{b_{\textrm{lim}}}\cap I_{{b_{\textrm{SC}}}}\geq\textrm{I}_{b_{\textrm{lim}}}$ implies that approximate SCC defines a tighter boundary than the grid-code value $\textrm{I}_{b_{\textrm{lim}}}$, therefore satisfying system security. Type-I errors are likely to cause insecure operating conditions, thus being undesired.
Evaluation 2: Minimum SCC distribution. We consider all buses to comply with approximate and linear SCC constraints respectively in an economic dispatch problem. The minimum SCC of all buses can then be calculated to verify if each of them at any time periods is beyond $\textrm{I}_{b_{\textrm{lim}}}$.
$\textrm{I}_{b_{\textrm{lim}}}$ (p.u.) Minimum SCC Type-I Type-II
Approximate (2a) Linearized (4c) $N_{e}$ errerr $N_{e}$ errerr
p.u. Bus p.u. Bus
3.00 4.50 30 4.73 30 0 0 2 -1.13%
4.00 4.13 30 4.06 30 0 0 8 -0.56%
5.00 5.04 26 5.20 30 0 0 11 -1.79%
Results evaluated at different $\textrm{I}_{b_{\textrm{lim}}}$ are presented in Table 2, in which $N_{e}$ and errerr respectively indicate total number and averaged value of approximation errors. The approximation eliminates Type-I errors and produces only a small amount of Type-II errors, indicating a good performance. The minimum SCC of the entire system is above the $\textrm{I}_{b_{\textrm{lim}}}$, meaning that the SCC level on all buses are always within the safe range, justifying the approximate SCC constraints (2a). The effectiveness of the linearized SCC constraint (4c) is demonstrated by the results on the middle section of Table 2, showing that minimum SCC is still beyond $\textrm{I}_{b_{\textrm{lim}}}$.
4.3 System Operation under Perfect Competition
The marginal SCC prices payable to different units for critical buses {26, 29, 30} are depicted in Fig. 3, in which the prices for bus 30 are in fact the same for several units (blue line in right plot) due to their very similar SCC contribution to this bus. For the case of bus 26 (left plot), only SGs in bus 27 would capture prices for SCC provision, whereas other SGs are not chosen to provide this service. This is due to the fact that 2gg-bb27 are closer to bus 26, providing a level of SCC which is already sufficient to secure that bus, while other units provide a lower volume, resulting in those SCC contributions being less cost-effective. This explains that the elimination of those excess contributions do not incur higher costs. Vice versa, without SCC provision from $g_{1}$-bb27 or $g_{2}$-bb27, another local unit or more distant units with limited SCC contribution would have to be committed for meeting the security requirement, rendering higher costs which lead to these marginal SCC prices for units in bus 27.
For bus 30 (right plot), more than half of SGs participate in this service market capturing a certain price, which is generally lower than that for the case of bus 26 since most units provide low SCC due to long electrical distances, implying that the service from any of them can be replaced by any other. The marginal service price for bus 29 is zero, meaning that it shows sufficient SCC once other buses are secured, so no extra units would be required.
4.4 System Operation under Imperfect Competition
The bilevel model introduced in Section 2.2 allows to represent the case where a power company may behave strategically in the market to increase profits. We assume this company to operate several SGs, so that market power issues may be observed through analyzing different cases:
Bus of strategic units
W Bus 2 Bus 3 Bus 4 Bus 5 Bus 27
1 3.60 423 4.07 341 6.79 155 7.24 146 7.39 292
[…]
100 -0.38 262 0.87 208 3.59 13 4.10 -3 2.78 103
1000 -0.38 261 0.87 197 3.59 12 4.10 -4 2.78 103
Table 4: Averaged Strategic Multipliers of SGs in the Energy Market
Strategic units Bus 2 Bus 3 Bus 4 Bus 5 Bus 27
$\sum_{t}\beta_{\hat{g},t}^{\textrm{m}}/T$ 1.71 1.67 1.54 1.29 1.05 1.03 1.00 1.00 1.04 1.00
[…]
Bus 2 Bus 3 Bus 4 Bus 5 Bus 27
[…]
4.4.1 Two Strategic SGs in a Single Bus
The strategic company in this setting owns two SGs placed in a single bus, thus producing five scenarios, i.e., buses {2, 3, 4, 5, 27} (bus 30 is not considered, as the units placed there are too expensive to be dispatched to provide services). The $r^{\textrm{DG}}$ (8) and profits (accounting for both energy and SCC markets) of the strategic company in all five cases are given in Table 3, with values for the penalty term W of {1, 10, 100, 1000}. As the penalty factor increases, $r^{\textrm{DG}}$ and the profits decrease. It is noteworthy that $r^{\textrm{DG}}$ could be negative, as the primal LL problem is relaxed to be linear, thus having a bound potentially lower than the dual [27]. From these results, we set going forward that the strategic firm gives equal weight to the DG (i.e., $\textrm{W}=\textrm{10}$) in all scenarios, since a noticeably small $r^{\textrm{DG}}$ can be achieved while yielding a substantial profit, the ultimate objective of the strategic company. Although higher profits could be obtained with the setting $\textrm{W}=\textrm{1}$, it actually suggests that the strategic firm can not make a reasonable evaluation regarding its profitability [25].
While market power from SCC services is the focus of this work, we still briefly analyze market power in the energy clearing, by checking averaged bidding decisions, i.e., $\sum_{t}\beta_{\hat{g},t}^{\textrm{m}}/T$, shown in Table 4. Units at buses {2, 3} are seen to strategically bid to maximize profits, whereas other expensive units normally offer true energy bids even if they are capable of misreporting them (note that a bidding multiplier of 1 indicates a truthful bid). This is because there are significantly cheaper units with important capacity in the energy market, making high-cost units difficult to be dispatched. Consequently, the optimal solution for them is bidding honestly in order to clear the market. Market power in SCC services is analyzed next.
[…]
4.4.2 Two Strategic SGs in Different Buses
[…]
5 Conclusion and Future Work
Given the issues identified with regards to imperfect competition in SCC markets, the main line for future work is to develop strategies to mitigate this market power. Potential countermeasures could be considered at both the market and technical levels, such as limiting the bid or clearing price of SCC services, weighing the negative impacts of installing new units in buses that are prone to trigger market power when planning the system, and investing in the nonprofitable synchronous condenser to a proper location.

### v3

@@ 3.2 Primal-Dual Single-Level Optimization Model
It is noteworthy that problem (9) is still nonlinear, as it still contains quadratic terms $\lambda^{\textrm{E}}_{t}P_{\hat{g},t}$, $\lambda_{b,t}^{\textrm{SCC}}\textrm{k}_{b\hat{g}}u_{\hat{g},t}$ in (4a), $\textrm{O}_{\hat{g},b,t}^{\textrm{SCC}}\beta_{\hat{g},b,t}^{\textrm{SCC}}u_{\hat{g},t}$, $\textrm{O}_{\hat{g}}^{\textrm{m}}\beta_{\hat{g},t}^{\textrm{m}}P_{\hat{g},t}$ in (5a). These bilinear terms can be linearized using McCormick envelopes with appropriate bounds on the associated variables, as demonstrated in Appendix A.
In summary, to solve the nonlinear bilevel model with non-convexity, binary variables are relaxed to be continuous to derive the dual of the primal LL problem, as seen in block (i) in Fig. 1. Then they are tightened to achieve realistic UC solutions, although this leads to a DG between the primal and dual LL problems. The penalty function method with parameter W is used to minimize the DG, giving a trade-off between the solution accuracy and values of objectives (block (ii)). Finally, nonlinear terms are restated to form the final single-level, mixed-integer linear program, in which the primal and dual LL variables are co-optimized with UL variables (block (iii)).
Figure 1: Diagram of solution methodology.
4 Case Studies
Here, the basic system setup is first described (Section 4.1) and the effectiveness of the SCC constraints is validated (Section 4.2). In the formal market power analysis, we begin by examining a perfectly competitive market case (Section 4.3), which serves as a basis for comparison with subsequent imperfectly competitive cases (Section 4.4). We then analyze the market outcomes under different settings in which the strategic player controls multiple SGs, including two strategic SGs located at the same node (Section 4.4.1) and cases with SGs located at different nodes (Section 4.4.2). In addition, we investigate possible measures to mitigate the adverse effects of market power (Section 4.5), and finally discuss how the adopted system and market settings may influence the analysis presented here (Section 4.6).
4.1 Test System Setting
Figure 2: Modified IEEE 30-bus system.
Table 1: Operating Parameters of Synchronous Generators
Bus 2 3 4 5 27 30
$\textrm{O}_{g}^{\textrm{nl}}$ (€/h) 1,743 1,501 1,376 1,093 990 857
$\textrm{O}_{g_{1}}^{\textrm{m}}$ (€/MWh) 6.20 7.10 10.47 12.28 13.53 15.36
$\textrm{O}_{g_{2}}^{\textrm{m}}$ (€/MWh) 7.07 8.72 11.49 12.84 14.60 15.02
$\textrm{K}_{g}^{\textrm{st}}$ (€/h) 20,000 12,500 9,250 7,200 5,500 3,100
$\textrm{K}_{g}^{\textrm{sh}}$ (€/h) 5,000 2,850 1,850 1,440 1,200 1,000
$\textrm{P}_{g}^{\textrm{min}}$ (MW) 658 576 302 133 130 58
$\textrm{P}_{g}^{\textrm{max}}$ (MW) 1,317 1,152 756 667 650 576
$\textrm{P}_{g}^{\textrm{rd}}$ (MW/h) 263 230 302 267 390 346
$\textrm{P}_{g}^{\textrm{ru}}$ (MW/h) 263 230 302 267 390 346
$u_{g,0}$ 1 1 1 1 1 0
$P_{g,0}$ (MW) 1,054 922 605 534 520 0
Table 2: Bus Locations of Generators
Generator type IBR SGs
   Bus 1, 23, 26 2, 3, 4, 5, 27, 30 (2 SGs/bus)
To examine whether market power issues may arise in SCC services, certain units under an appropriate system scale should be selected, a modified IEEE 30-bus system (as seen in Fig. 2) is therefore adopted in the case studies. For other given systems, such as larger-scale ones, the proposed methodology can be applied to analyze such concerns if they indeed exist. The IBR, i.e., wind turbines are added to buses {1, 23, 26}, with energy bids {2.46, 3.19, 2.86} €/MWh, respectively. In order to model a modular SCC contribution from SGs, we consider that two units are located in each bus containing an SG, i.e., buses {2, 3, 4, 5, 27, 30}. This way, diverse decisions can be made in the market. For example, different UC status (i.e., both online, both offline, just one of them online) can determine various SCC contributions and bids. The SGs in bus 2 are separately named as $g_{1}$-bb2 and $g_{2}$-bb2, with 2gg-bb2 standing for both of them. Other thermal units follow the same nomenclature. The SGs’ parameters are given in Table 1. The locations of units are summarized in Table 2.
The network impedances and reactances of SGs refer to [29] and [30], respectively. Load demand $\textrm{P}^{\textrm{D}}_{t}\in[\textrm{5109},\textrm{7689}]$ MW. Total installed wind generation capacity is 750 MW. Base power is $S_{B}=\textrm{100}~\textrm{MVA}$. $\textrm{I}_{c}$, the short-circuit injection from IBR, is 1.00 p.u. Nominal voltage is set as 0.95 p.u. for conservativeness [8]. Parameter $\nu$ is set as 0.1 to yield a desired approximation for SCC constraints. Maximum multipliers for bidding decisions $\overline{\upbeta}^{\textrm{m}}_{\hat{g}}$ and $\overline{\upbeta}^{\textrm{SCC}}_{\hat{g},b}$ are both set as 2. All code and data are available at the public repository [27]. The simulations in this work were processed by Julia-JuMP and Gurobi in version 12.0.0. The Gurobi solver is employed with default settings, where the key parameters are set as follows: the relative gap tolerance (MIPGap) is 1e-4, the feasibility tolerance (FeasibilityTol) and optimality tolerance (OptimalityTol) are both 1e-6, and the integer feasibility tolerance (IntFeasTol) is 1e-5. Other solver parameters (e.g., threads, presolve) use the solver’s auto-selected defaults.
4.2 Validation of Short-Circuit Current Constraints
The approximate SCC constraints (2a) are linearized to be (5c) for the primal-dual formulation. Two evaluations are first conducted to validate the obtained SCC constraints.
Evaluation 1: Training errors of approximation. Two types of errors are analyzed to verify the approximation [7]: Type-I: $I_{b_{\textrm{L}}}\geq\textrm{I}_{b_{\textrm{lim}}}\cap I_{{b_{\textrm{SC}}}}<\textrm{I}_{b_{\textrm{lim}}}$ means the approximate SCC is looser than the actual requirement, a potentially dangerous situation; Type-II: $I_{b_{\textrm{L}}}<\textrm{I}_{b_{\textrm{lim}}}\cap I_{{b_{\textrm{SC}}}}\geq\textrm{I}_{b_{\textrm{lim}}}$ implies that the approximate SCC defines a tighter boundary than the grid-code value $\textrm{I}_{b_{\textrm{lim}}}$, therefore satisfying system security. Type-I errors are likely to cause insecure operating conditions, thus being undesired. The averaged value of the errors is defined below:
$N_{e}=|\xi|,~err=\frac{1}{N_{e}}\sum_{\xi}\frac{I^{(\xi)}_{b_{\textrm{L}}}-I^{(\xi)}_{b_{\textrm{SC}}}}{I^{(\xi)}_{b_{\textrm{SC}}}}$ (11)
where $\xi$ indicates the enumerated system operating points. $N_{e}$ is the total number of errors.
Evaluation 2: Minimum SCC distribution. We consider all buses to comply with approximate and linear SCC constraints respectively in an economic dispatch problem. The minimum SCC of all buses can then be calculated to verify whether the SCC at each bus and at each time period remains above $\textrm{I}_{b_{\textrm{lim}}}$.
$\textrm{I}_{b_{\textrm{lim}}}$ (p.u.) Minimum SCC Type-I Type-II
Approximate (2a) Linearized (5c) $N_{e}$ errerr $N_{e}$ errerr
p.u. Bus p.u. Bus
3.00 4.50 30 4.73 30 0 0 2 -1.13%
4.00 4.13 30 4.06 30 0 0 8 -0.56%
5.00 5.04 26 5.20 30 0 0 11 -1.79%
Results evaluated at different $\textrm{I}_{b_{\textrm{lim}}}$ are presented in Table 3. It is shown that the approximation eliminates Type-I errors and produces only a small amount of Type-II errors, indicating a good performance. The minimum SCC of the entire system is above the $\textrm{I}_{b_{\textrm{lim}}}$, meaning that the SCC level on all buses is always within the safe range, justifying the approximate SCC constraints (2a). The effectiveness of the linearized SCC constraint (5c) is demonstrated by the results on the middle section of Table 3, showing that minimum SCC is still beyond $\textrm{I}_{b_{\textrm{lim}}}$.
4.3 System Operation under Perfect Competition
Table 4: SCC Contributions from SGs to Critical Buses (p.u.)
SGs contributing SCC SCC captured by critical buses
Bus 26 Bus 29 Bus 30
$g_{1}$-$g_{2}-bb2$ 2.75/2.75 2.95/2.95 2.91/2.92
$g_{1}$-$g_{2}-bb3$ 2.76/2.77 2.96/2.96 2.91/2.91
$g_{1}$-$g_{2}-bb4$ 2.77/2.77 2.97/2.99 2.91/2.91
$g_{1}$-$g_{2}-bb5$ 2.76/2.77 2.97/2.99 2.92/2.92
[…]
The marginal SCC prices payable to different units for critical buses {26, 29, 30} are depicted in Fig. 4, in which the prices for bus 30 are in fact the same for several units (blue line in the lower plot) due to their very similar SCC contribution to this bus. For the case of bus 26 (upper plot), only SGs in bus 27 would capture prices for SCC provision, whereas other SGs are not chosen to provide this service. This is due to the fact that 2gg-bb27 are closer to bus 26, providing a higher level of SCC which is already sufficient to secure that bus (as seen in Table 4), while other units provide a lower volume, resulting in those SCC contributions being less cost-effective. This explains that the elimination of those excess contributions does not incur higher costs. Vice versa, without SCC provision from $g_{1}$-bb27 or $g_{2}$-bb27, another local unit or more distant units with limited SCC contribution would have to be committed for meeting the security requirement, rendering higher costs which lead to these marginal SCC prices for units in bus 27.
For bus 30 (lower plot), more than half of SGs participate in this service market capturing a certain price, which is generally lower than that for the case of bus 26 since most units provide low SCC due to long electrical distances, implying that the service from any of them can be replaced by any other. The marginal service price for bus 29 is zero, meaning that it shows sufficient SCC once other buses are secured, so no extra units would be required.
4.4 System Operation under Imperfect Competition
The bilevel model introduced in Section 2.2 allows to represent the case where a power company may behave strategically in the market to increase profits. We assume this company to operate several SGs, so that market power issues may be observed through analyzing different cases.
Bus of strategic units
W Bus 2 Bus 3 Bus 4 Bus 5 Bus 27
1 3.60 423 4.07 341 6.79 155 7.24 146 7.39 292
[…]
100 -0.38 262 0.87 208 3.59 13 4.10 -3 2.78 103
1000 -0.38 261 0.87 197 3.59 12 4.10 -4 2.78 103
Table 6: Averaged Strategic Multipliers of SGs in the Energy Market
Strategic units Bus 2 Bus 3 Bus 4 Bus 5 Bus 27
[…]
Bus 2 Bus 3 Bus 4 Bus 5 Bus 27
[…]
4.4.1 Two Strategic SGs in a Single Bus
The strategic company in this setting owns two SGs placed in a single bus, thus producing five scenarios, i.e., buses {2, 3, 4, 5, 27} (bus 30 is not considered, as the units placed there are too expensive to be dispatched to provide services). The $r^{\textrm{DG}}$ (10) and profits (accounting for both energy and SCC markets) of the strategic company in all five cases are given in Table 5, with values for the penalty term W of {1, 10, 100, 1000}. As the penalty factor increases, $r^{\textrm{DG}}$ and the profits decrease. It is noteworthy that $r^{\textrm{DG}}$ could be negative, as the primal LL problem is relaxed to be linear, thus having a bound potentially lower than the dual [28]. From these results, we set going forward that the strategic firm gives equal weight to the DG (i.e., $\textrm{W}=\textrm{10}$) in all scenarios, since a noticeably small $r^{\textrm{DG}}$ can be achieved while yielding a substantial profit, the ultimate objective of the strategic company. It can also ensure that the weight between objectives in both levels are the same in all scenarios. Although higher profits could be obtained with the setting $\textrm{W}=\textrm{1}$, it actually suggests that the strategic firm can not make a reasonable evaluation regarding its profitability [26]. While a larger W can only slightly reduce the gap, it noticeably lowers the profit at the upper level, thereby violating the criterion for selecting W. A comparable parameter selection strategy for balancing the objectives in (9a) can be found in [9], where the authors achieved similar performance in solving the bilevel problem, guaranteeing desirable profits for strategic agents with acceptable market clearing accuracy.
While market power from SCC services is the focus of this work, we still briefly analyze market power in the energy clearing, by checking averaged bidding decisions, i.e., $\sum_{t}\beta_{\hat{g},t}^{\textrm{m}}/T$, shown in Table 6. Units at buses {2, 3} are seen to strategically bid to maximize profits, whereas other expensive units normally offer true energy bids even if they are capable of misreporting them (as illustrated by the shaded area in comparison with other strategic scenarios; note that a bidding multiplier of 1 indicates a truthful bid). This is because there are significantly cheaper units with important capacity in the energy market, making high-cost units difficult to be dispatched. Consequently, the optimal solution for them is bidding honestly in order to clear the market. Market power in SCC services is analyzed next.
[…]
4.4.2 Two Strategic SGs in Different Buses
[…]
4.5 Potential Countermeasures to Mitigate Market Power
Countermeasures to limit market power in SCC services can be broadly categorized into two types: market and technology-based.
4.5.1 Market-Based Measures
Upper limit of secure SCC: An upper bound on SCC constraints can be imposed at critical buses to enable greater participation of units from different locations in the market, thereby maintaining SCC within a limited range. For example, setting an appropriate SCC limit at bus 26 could enforce SGs at bus 27, which contribute significantly to SCC, to be shut down, while other units with smaller SCC contributions are dispatched as substitutes. However, determining an appropriate upper bound remains challenging. If the limit is set too low, essential SGs may be forced offline, especially given that SCC is provided in a discrete manner. Further analysis on this approach is necessary to conclude its suitability for a given system.
Service price cap: Setting a cap on SCC service prices could directly limit market power. However, determining an appropriate cap would require accurate evaluation of service offers submitted by market participants to ensure efficient market clearing. In this study, the bidding multiplier is capped at 2, which is adopted as a conservative upper-bound assumption to fully characterize the maximum potential of strategic SGs to exercise market power. A key issue is that, in practice, the bidding decision does not necessarily reflect the true marginal cost perceived by the service provider. Even with a relatively low bid, an agent may claim a higher marginal service cost; conversely, a higher bid may be associated with a lower actual cost. Ultimately, system operators can only observe the submitted service offers. Therefore, a critical challenge is that some power companies may somehow misreport their offers, which makes it difficult for regulators or system operators to achieve cost-effective system operation [31].
Demand response: This market mechanism allows consumers to flexibly curtail or shift their loads when SCC service prices rise or when SGs with market power operate for extended periods. Such flexible load adjustments not only have the potential to restrain the strategic commitment behavior of these SGs (since lower loading conditions may cause certain units to go offline), but also enable consumers to receive economic compensation for their load regulation efforts, thereby directly alleviating the adverse economic impacts of strategic SGs’ behavior on the demand side. A preliminary analysis on the role of demand response in SCC-constrained UC has been conducted in [32], laying the foundation for further investigating its impact on the market power of SCC services.
4.5.2 Technology-Based Measures
SCC support devices: A feasible solution is to deploy local SCC response devices or other auxiliary technologies, such as synchronous condensers. It is evident that if a synchronous condenser is installed on bus 26, the market power in this paper will be mitigated significantly as the SCC level for that bus could be satisfied without requiring further SCC services. However, it is relevant to analyze who should own and be responsible for the investment cost in these devices.
System topology: Analyzing the electrical topology of the system to assess short-circuit risks at each bus and to identify locations that may give rise to market power could also be an effective approach. In this way, investment decisions can better balance system benefits against the number of units installed at such buses, and the network topology could also be reconfigured to mitigate market power. The latter approach would entail additional complexity for system operators, which should be further analyzed.
4.6 Potential Impact of Model Settings on Market Power in SCC Services
In this subsection, we clarify the potential impact of system assumptions in this paper on the manifestation of market power for SCC.
4.6.1 Uniform SCC Threshold
In the present work, a uniform secure SCC threshold, $\textrm{I}_{b_{\textrm{lim}}}$, is adopted for all buses. In practice, however, the minimum acceptable SCC level may vary depending on the bus type due to factors such as protection configurations, and it should be confirmed before the system scheduling. This work is not intended to determine $\textrm{I}_{b_{\textrm{lim}}}$ for each individual bus, but to propose a market framework that enables relevant agents to examine market power using given SCC thresholds. In order to provide a practical and conservative SCC security, $\textrm{I}_{b_{\textrm{lim}}}=5~\textrm{p.u.}$ is adopted in the case studies, which guarantees strong grid conditions at the buses of interest, as specified in [8].
Furthermore, this assumption does not undermine the results of this study, as the local or remote coupling between SGs and buses (the fundamental driver of market power) is inherently dependent on the electrical distance, rather than on $\textrm{I}_{b_{\textrm{lim}}}$. It is worth noting, however, that this setting may instead lead to an underestimation of market power for SGs located near buses with stricter practical SCC requirements (e.g., grid-following IBR buses that may require a higher level of SCC). Such buses rely more on electrically proximate SGs, resulting in lower service substitutability. In the system studied in this paper, these factors would amplify the market power of SGs at bus 27 with respect to bus 26, underscoring the importance of carefully designing the SCC ancillary service market.
4.6.2 Approximated SCC Representation
Another point worth noting is that the approximated SCC levels (2a) may introduce deviations in market outcomes, such as strategic behaviors and SCC clearing prices. Nevertheless, this bias remains within a controllable range and affects system-wide SCC levels rather than those at specific buses, implying that the core conclusion that market power is fundamentally driven by system topology and location of the different assets would not be altered.
From a system stability perspective, this approximation error is not of concern, as the condition $\textrm{I}_{b_{\textrm{lim}}}\leq I_{b_{\textrm{L}}}\leq I_{b_{\textrm{SC}}}$ holds, indicating that the classification inherent in the approximation process introduces an implicit safety margin into the system. This observation is further supported by the training performance of the Type-II error computed by (11) and reported in the last column of Table 3, where the mean value of this error remains negative.
4.6.3 Uncertainty of Wind Generation
The impact of wind power capacity on market power in SCC services primarily arises through its coupling with the commitment of SGs. This is because, on the one hand, wind-driven IBR contribute significantly less to SCC than SGs and would be therefore insufficient for themselves to exert market power; on the other hand, the net load (system energy demand net of available wind power) directly affects the commitment status of SGs, and consequently the supply of SCC.
A phenomenon one can expect is that, under high net-load conditions, SGs are committed more frequently to satisfy energy demand and may simultaneously exert market power in SCC. By contrast, when load demand is low or wind power availability is high, SGs with lower costs and appropriate SCC capacity are more likely to exert market power, as they are preferentially dispatched by the system operator, while other higher-cost units may not be dispatched at all.
Consequently, the uncertainty of wind power output affects market power in SCC services in an indirect manner. Specifically, as long as wind generation remains at a level that does not alter the commitment status of SGs, its uncertainty would not, in theory, significantly change market outcomes. However, when available wind resources are close to the level at which certain SGs need to switch their operating states, such uncertainty may substantially mitigate or exacerbate market power, depending on whether the SGs exerting market power are switched to offline or online.
4.6.4 Single-Node Energy Price
The UC model in this paper adopts the simplifying assumption of single-node electricity price. When congestion exists, the spatial heterogeneity of Locational Marginal Prices (LMPs) is coupled with the locational characteristics of SCC prices; however, this coupling would be implicit. This is because the SCC price is directly related to binary unit commitment variables, whereas congestion is capacity-related and only indirectly linked to unit commitment decisions. Future research should explicitly characterize this relationship.
4.6.5 Single-Day System Scheduling
For longer multi-day scheduling horizons with more diverse system operating conditions, actual generating units will optimize their cross-period schedules. This may either mitigate or enhance the short-term revenue gains of strategic agents, depending on the actual electricity demand and available wind power, as discussed in Sections 4.5.1 and 4.6.3, but cannot eliminate the inherent market power arising from the grid topological heterogeneity in SCC provision.
5 Conclusion and Future Work
Detailed market power studies for specific power systems should be carried out before determining if a market for SCC is suitable for those systems, given that alternative SCC procurement mechanisms (e.g., cost-based contracts, grid codes) would also be feasible complements for bus-specific SCC needs. Given the issues identified with regards to imperfect competition in SCC markets, the main line for future work is to develop strategies to mitigate this market power. Potential countermeasures could be considered at both the market and technology levels, such as limiting the clearing price of SCC services, weighing the negative impacts of installing new units in buses that are prone to trigger market power when planning the system. Investing in synchronous condensers at proper locations would also be effective, making it relevant to analyze who should own and bear the investment costs. Moreover, market participants that may influence SCC provision, such as load aggregators, would be incorporated into future research to further refine the analysis of strategic behavior in SCC ancillary service markets.
Appendix A McCormick Envelopes
The McCormick envelopes with appropriate bounds, as suggested in [26], are constructed for the relevant bilinear terms, as presented below.
@@ Appendix A McCormick Envelopes
Acknowledgment
This work was supported by MICIU/AEI/10.13039/501100011033 and ERDF/EU under grant PID2023-150401OA-C22, as well as by the Madrid Government (Comunidad de Madrid-Spain) under the Multiannual Agreement 2023-2026 with Universidad Politécnica de Madrid, ‘Line A - Emerging PIs’ (grant number: 24-DWGG5L-33-SMHGZ1). The work of Peng Wang was also supported by China Scholarship Council under grant 202408500065.